MC algo 11-7 Portfolio Expected Return You have gathered the following information on your investments.

What is the expected return on the portfolio?

Stock # of Shares Price per Share Expected Return

F 390 $ 48 13.64 %

G 355 $ 34 10.35 %

H 295 $ 60 10.83 %

Multiple Choice:

A. 11.61%

B. 13.18%

C. 12.72%

D. 11.80%

E. 12.26%
Answer #1

Expected return on portfolio = (wF × rF) + (wg+ rG) + (wH × rH)

W= weight = Stock price / total portion of value

Total portion of value = share price F + share price G + share price H

Total portion of value = (390×48) +(355× 34) + (295×60)

Total portion of value= 18960+12070+17700= 48730

r= rate of return

Weight of F= 18960/48730=0.389083

Weight of G= 12070/48730= 0.247691

Weight of H= 17700/48730= 0.363226

Expected return on portfolio= (0.389083×0.1364) + (0.247691×0.1035) + (0.363226×0.1083)

Expected return on portfolio=(0.053071+0.025636+0.039337)

Expected return on portfolio = 0.118044 or 11.80%

D is correct answer